OpenStudy (anonymous):

What is the value of y in the solution to the following system of equations? 2x + y = −4 5x + 3y = −6

OpenStudy (anonymous):

x = -6, y = 8

OpenStudy (anonymous):

if you have a texas instrument, its really easy to solve these. put them in a matrix. (2nd + x^-1 buttons), then scroll over to "edit" at the top, enter 2 x 3, then type in all your values which would look like this 2 1 -4 5 3 -6 then hit (2nd + mode) to exit. then go back to the matrix (2nd + x^-1) buttons and go over to math. press the option B, which is "rref(" and click enter. it'll enter that value, so go back to (2nd + x^-1) and select your matrix from that screen and hit enter. your new matrix will look like this: 1 0 -6 0 1 8 1 0; 0 1 are your multiplication values, just ignore those. -6 is your value for x, and 8 your value for y.

OpenStudy (texaschic101):

Substitution method : 2x + y = - 4 --> y = -2x - 4 now sub -2x - 4 in for y in the other equation 5x + 3y = - 6 5x + 3(-2x - 4) = - 6 5x - 6x - 12 = - 6 -x = - 6 + 12 -x = 6 x = - 6 now sub -6 in for x in either of the original equations 2x + y = - 4 2(-6) + y = - 4 - 12 + y = - 4 y = -4 + 12 y = 8 check... 2x + y = - 4 2(-6) + 8 = - 4 -12 + 8 = - 4 -4 = -4 (correct) ANSWER : x = -6 and y = -4